# Enhancing Portfolio Optimization with Transformer-GAN Integration: A Novel Approach in the Black-Litterman Framework

### v2

@@ I Introduction
The Black-Litterman model (BL), a renowned tool in modern portfolio theory, offers a structured method for combining investor views with market equilibrium, thereby enabling more tailored and robust portfolio allocations. It significantly mitigates the limitations inherent in the traditional Markowitz mean-variance optimization, especially those related to input sensitivity and estimation errors[1][2]. However, generating subjective views, a critical component of this model, remains a challenge.
With the advancement of artificial intelligence techniques, scholars have explored various methodologies to generate investor views for the Black-Litterman (BL) model. CNNs, by leveraging multiple indicators, and LSTMs, through historical price data analysis, have significantly contributed to enhancing portfolio performance within the BL framework[3][4][5]. However, the complexity of processing numerous indicators increases computational demands, while LSTMs face challenges in capturing long-range dependencies[6]. This necessitates the exploration of new models that can address these limitations and improve price prediction accuracy.
Originally designed for language translation, Transformers have demonstrated exceptional versatility across domains such as computer vision[7], speech recognition[8], and Large Language Models (LLMs)[6]. Their unique architecture, featuring attention mechanisms, enables effective sequential data management, surpassing traditional recurrent and convolutional network models. In the realm of stock price prediction, studies have indicated that Transformer-based models outperform their CNN and LSTM counterparts[9][10].
Furthermore, Generative Adversarial Networks (GANs)[11] which are constructed by generator and discriminator, have proven adept at uncovering complex patterns in historical stock data, a task challenging for simpler models. Hence, GAN-based methodologies excel in time series forecasting and stock prediction, offering innovative performance improvements[12][13].
In the pursuit of enhancing the Black-Litterman model’s capabilities, we have innovated an architecture that synergizes the generative power of GANs with the intricate data processing prowess of Transformers. The resulting framework is aimed at capturing the intricacies and potential patterns within financial time series data, which can be used to generate objective views for Black-Litterman to enhance portfolio construction. Here are the salient features of our approach:
@@
  • Advanced Architectural Integration: Our model innovatively combines the generative capabilities of GANs with the deep learning prowess of Transformers, adept at uncovering long-range dependencies and latent patterns within historical stock price data.
  • Performance and Adaptability: The proposed model demonstrates competitive performance metrics, establishing new frontiers in accuracy and scalability, which are corroborated by rigorous benchmarking against traditional GANs and showcasing exceptional adaptability across various data resolutions.
@@
  • Strategic Asset Allocation: The application of our integrated model to the Black-Litterman framework facilitates a more nuanced and adaptive approach to portfolio construction. By accurately capturing market trends and investor sentiments, it guides asset allocation decisions that can potentially outperform standard market strategies.
@@ II Related work
The application of the Black-Litterman (BL) model is contingent upon the precise estimation of parameters derived from return series data. This section focuses on exploring mechanisms that bolster the efficacy of historical data in enhancing portfolio performance. Thus, we scrutinize cutting-edge deep learning algorithms renowned for their predictive prowess and exceptional capability in capturing latent patterns and salient features within financial datasets.
@@
Convolutional Neural Networks (CNNs) utilize convolutional kernels to autonomously extract pivotal features from stock price charts or technical indicators, thereby achieving remarkable accuracy [14] [15]. Yet, CNNs do not inherently account for the sequential nature of stock prices, necessitating additional transformations of time-series data [16]. Addressing this, Long Short-Term Memory networks (LSTM), a specialized subset of Recurrent Neural Networks (RNN), employ ‘forget gate’ and ‘input gate’ mechanisms to modulate information flow, significantly boosting their efficacy in processing time-series data in stock analysis [14] [17]. However, despite their prowess in temporal feature capture, LSTMs are less adept at extracting features from high-dimensional spatial data. To mitigate the shortcomings of singular models in predicting stock prices, the CNN-LSTM hybrid model has been proposed, merging CNNs’ robust feature extraction capabilities with LSTMs’ proficiency in time-series analysis [16] [18] [19]. Further advancements include integrating the attention mechanism into the CNN-LSTM framework, sharpening the model’s focus on crucial temporal information and thereby enhancing prediction accuracy and efficiency [20] [21] [22]. Leveraging their proficiency with time series data, models like CNN, LSTM, and their hybrids have been instrumental in forecasting future prices and integrating these predictions as views into the Black-Litterman framework. Barua et al.[5] introduced a CNN-BiLSTM model for predicting index closing prices, subsequently incorporating these forecasts as investor views within the Black-Litterman model. Similarly, Rezaei et al.[4] demonstrated that the hybrid CEEMD-CNN-LSTM model achieves outstanding performance in view generation, highlighting the potential of combining advanced machine learning techniques for financial forecasting and portfolio optimization.
The attention mechanism[23], pivotal to the transformer model, has revolutionized sequential data analysis, enhancing models’ focus on pertinent input segments for more effective processing, particularly with lengthy sequences. The transformer distinguishes itself by processing data in parallel, its scalability, and handling long-range dependencies, significantly improving prediction accuracy and modeling complex patterns. While Chen et al.[24] and Zhang et al.[25] have straightforwardly utilized the Transformer for stock market forecasts, its self-attention mechanism’s ability to capture long-term dependencies is particularly notable. This feature has led Wang et al.[9] and Zhang et al.[10] to successfully employ the transformer for stock price predictions, markedly outperforming traditional methods. The model’s adaptability to various data lengths and parallel processing capabilities make it a powerful tool for detailed and accurate large-scale data analysis.
Generative Adversarial Networks (GANs)[11] have shown remarkable capability in capturing potential patterns or features from complex datasets for their potential to learn high-dimensional, complex data distribution[26]. For instance, Zhang et al. discuss the use of GANs in recovering hidden cell-lineage developmental knowledge from complex biological datasets, showcasing the model’s ability to capture essential data semantics in its latent space [27]. Moreover, GANs’ application in anomaly detection further exemplifies their capacity to identify abnormal instances by learning potential patterns within data, as explored by Rani et al.[28]. GANs have shown promise in the domain of time series forecasting. Ferchichi et al.[29] introduced a novel multi-attention GAN model, comprising three main components, which achieves notable accuracy across multiple future time intervals. Yin et al.[30] employed a multi-attention GAN framework for stock market prediction, while Polamuri et al.[13] introduced GAN-HPA, yielding promising results in stock price forecasting. Similarly, study of Diqi et al.[12] on StockGAN showcases its efficacy in financial time series analysis, underscoring the growing utility of GAN-based models in predicting market dynamics.
@@ III-A GAN
The realm of Artificial Intelligence (AI) has seen remarkable advancements, leading to significant enhancements in machines’ capabilities to mimic human traits. A central strategy in emulating authentic human behavior is through the development of sophisticated machine learning algorithms. Among these, Generative Adversarial Networks (GANs), introduced by Goodfellow et al. in 2014[11], stand out for their unique operational framework. GANs are predicated on the concept of a two-player zero-sum game, where the sum of the gains and losses between the players is zero, ensuring that any gain or loss in utility by one player is exactly offset by the loss or gain in the other player’s utility.
@@
III-B Transformer
Radford et al.[6] conducted ablation studies to assess the Transformer architecture’s effectiveness in comparison to models based on Long Short-Term Memory (LSTM). They observed a significant reduction in performance metrics, such as accuracy and training efficiency, when employing LSTM models. This decline is attributed to LSTMs’ inherent limitation in capturing long-distance dependencies due to issues like the vanishing gradient problem, restricting their ability to model complex linguistic details effectively. Additionally, CNN has the same limitation of processing long-term temporal dependencies[31]. In contrast, The adoption of Transformer networks, with their self-attention mechanism, markedly improves the model’s capacity to understand extended linguistic structures[32][33][34]. This advantage stems from the Transformers’ ability to weigh the importance of all parts of the input data, regardless of their position, facilitating a deeper comprehension of language and context-based content. The key principle of transformer is attention mechanism.
Classical attention mechanisms, such as traditional, multi-head, and temporal attention, play a pivotal role across a broad spectrum of applications. These mechanisms are instrumental in advancing fields like text classification[6], speech recognition[8], image recognition[7], and video processing[35][36]. Each type of attention mechanism offers unique advantages in processing and interpreting complex data structures, thereby contributing significantly to improvements in these areas.
In this study, we have employed the aforementioned attention mechanisms to adeptly capture contextual nuances and salient features inherent in historical stock price data.
@@ III-C Black-Litterman Framework
The Mean-Variance (MV) model, while foundational in portfolio optimization, has limitations, particularly its sensitivity to input estimates, which can lead to non-robust portfolio allocations. This sensitivity can result in portfolios that are not truly diversified and may overconcentrate in a few assets[37] Moreover, the MV model does not adequately address the asymmetry of stock returns and can underestimate risk during market downturns. These shortcomings justify the need for more robust approaches, such as the Black-Litterman model, which integrates market equilibrium with investor views to enhance portfolio construction[38], aiming for more intuitive and stable asset allocation[39], addressing the MV model’s issues by generating portfolios with lower risk, less extreme asset allocations, and higher diversification across asset classes [40].
This equation represents the calculation of the equilibrium return vector, $\Pi$, where $\delta$ denotes the risk aversion coefficient, $\Sigma$ is the covariance matrix of asset returns, and $w_{marker}$ are the market capital weights.
@@ IV-A2 The Superior Features of Transformer Models
In the context of transformer models, this is typically realized through the scaled dot-product attention, which can be mathematically expressed as:
Here, $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively, derived from the input data, and $d_{k}$ is the dimension of the key vector. This formula allows the model to dynamically assign importance to different parts of the input sequence based on their relevance to the current processing context.
@@
The attention function is then applied in parallel, yielding multiple output values that are concatenated and once again linearly projected. This mechanism is represented as:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},...,\text{head}_{h})W^{O}$ (11)
where each head is defined as:
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (12)
Here, $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, and $W^{O}$ are parameter matrices, and $h$ is the number of heads.
@@ Acknowledgment
The preferred spelling of the word “acknowledgment” in America is without an “e” after the “g”. Avoid the stilted expression “one of us (R. B. G.) thanks $\ldots$”. Instead, try “R. B. G. $\ldots$”. Put sponsor acknowledgments in the unnumbered footnote on the first page.

### v3

@@ I Introduction
The Black-Litterman (BL) model, foundational in modern portfolio theory, adeptly merges market equilibrium with investor insights to refine portfolio allocations. This integration addresses key limitations of the traditional Markowitz mean-variance optimization, particularly issues related to input sensitivity and estimation errors[1][2]. The generation of accurate and objective investor views remains a significant challenge, underlining the necessity for advanced, more effective methods.
Recent advancements in artificial intelligence provide promising solutions for generating these views. CNNs and LSTMs have been instrumental within the BL framework, leveraging a range of data indicators and historical price analyses to enhance portfolio performance[3][4][5]. However, their effectiveness is often hampered by the complexity of processing multiple indicators and their struggles with long-term dependencies[6]. These challenges highlight the need for innovative and robust approaches that can better capture and utilize the intricacies of financial data.
Enter the Transformer and Generative Adversarial Network (GAN) models, which represent the cutting edge of deep learning technology. Originally developed for natural language processing, Transformers have shown remarkable versatility and superior performance in various domains, including computer vision[7] and speech recognition[8], due to their unique architecture that facilitates effective sequential data handling[6]. Similarly, GANs have emerged as powerful tools for modeling and generating complex data distributions, proving particularly effective in unveiling latent patterns within historical financial data[9].
This study introduces a novel integration of GANs with Transformer models within the Black-Litterman framework to optimize portfolio construction. This hybrid approach not only enhances the generation of investor views by distilling complex financial time series data into actionable insights but also significantly improves the predictive accuracy and robustness of portfolio allocations. Our contributions can be summarized in three main advancements:
@@
  • Advanced Architectural Integration: Our model uniquely combines the generative capabilities of GANs with the deep learning efficiency of Transformers, making it particularly adept at uncovering long-range dependencies and latent patterns within historical stock price data.
  • Performance and Adaptability: Through rigorous benchmarking, our model demonstrates superior accuracy and scalability, adapting seamlessly to various data resolutions and market conditions.
@@
  • Strategic Asset Allocation: The integration of our model with the Black-Litterman framework allows for a more nuanced and adaptive approach to portfolio construction, significantly enhancing the potential to outperform traditional market strategies.
@@ II Related Work
II-A Enhancements in Portfolio Optimization Models
The Mean-Variance (MV) model, foundational in portfolio optimization, has notable limitations, particularly its sensitivity to input estimates which can lead to non-robust portfolio allocations [10]. These portfolios often lack true diversification and may concentrate heavily in a few assets. Moreover, the MV model struggles with the asymmetry of stock returns and often underestimates risk during market downturns. To address these issues, more robust models such as the Black-Litterman (BL) model have been developed. The BL model integrates market equilibrium with investor views to enhance portfolio construction [11], aiming for more intuitive and stable asset allocation [12]. This model addresses the limitations of the MV model by generating portfolios with lower risk, less extreme asset allocations, and higher diversification across asset classes [13].
II-B Deep Learning Approaches in Financial Markets
@@
Deep learning algorithms, particularly Convolutional Neural Networks (CNNs), utilize convolutional kernels to autonomously extract pivotal features from stock price charts or technical indicators, achieving remarkable accuracy [14] [15]. However, CNNs do not inherently account for the sequential nature of stock prices, necessitating additional transformations of time-series data [16]. Addressing this, Long Short-Term Memory networks (LSTM), a specialized subset of Recurrent Neural Networks (RNN), employ mechanisms to modulate information flow, significantly boosting their efficacy in processing time-series data in stock analysis [14] [17]. Despite their prowess in temporal feature capture, LSTMs are less adept at extracting features from high-dimensional spatial data. The CNN-LSTM hybrid model has been proposed to merge CNNs’ robust feature extraction capabilities with LSTMs’ proficiency in time-series analysis [16] [18] [19]. Further advancements include integrating the attention mechanism into the CNN-LSTM framework, enhancing prediction accuracy and efficiency [20] [21] [22].
II-C Transformer and Attention Mechanisms
The attention mechanism, pivotal to the Transformer model, has revolutionized the field of sequential data analysis by enhancing the model’s focus on relevant segments of input data, particularly beneficial for processing lengthy sequences [23]. This innovation allows Transformers to process data in parallel, handle long-range dependencies, and scale efficiently, enhancing prediction accuracy and the modeling of complex patterns. Its effectiveness is demonstrated in stock market forecasting, where its self-attention mechanism captures long-term dependencies [24] [25]. In comparison, LSTM models, due to their limited ability in capturing long-distance dependencies, often fall short in performance metrics such as accuracy and training efficiency [6]. This limitation is also noted in CNN models, which struggle with processing long-term temporal dependencies [26]. The Transformers’ self-attention mechanism effectively weighs the importance of all parts of the input data, irrespective of their position [27] [28] [29]. Sun et al. employ a modified Transformer network, stripped of the position encoding module, to learn the policy for determining the subjective views of expected return for the Black-Litterman model [30].
II-D Advancements in Generative Models for Financial Analysis
Generative Adversarial Networks (GANs) stand out for their unique operational framework, predicated on the concept of a two-player zero-sum game [9]. These models have shown remarkable capability in capturing potential patterns or features from complex datasets [31]. For instance, GANs have been used in anomaly detection, showcasing their capacity to identify abnormal instances by learning potential patterns within data [32]. GANs have also shown promise in the domain of time series forecasting, with models like StockGAN demonstrating efficacy in financial time series analysis [33].
@@ III-A GAN
Generative Adversarial Networks (GANs) feature two neural models—the Generator (G) and the Discriminator (D)—in a competitive setup that produces high-quality synthetic data closely mimicking real distributions. This section explains their roles and the mathematics underlying their interaction.
@@
III-B Transformer Mechanism
The Transformer architecture, introduced by Vaswani et al.[23], has revolutionized the field of deep learning by offering an alternative to recurrent neural networks (RNNs) for processing sequential data. Central to the Transformer’s architecture is the self-attention mechanism, which enables the model to weigh the significance of different parts of the input data without the sequential processing limitations of RNNs.
The key innovation of the Transformer is its ability to handle dependencies between any two elements in the input data, regardless of their positions. This capability is achieved through the self-attention mechanism, which computes a weighted sum of all input elements, with weights assigned based on a function of the elements’ compatibility. The mathematical representation of the self-attention mechanism is given by:
where $Q$, $K$, and $V$ represent the queries, keys, and values matrices respectively, and $d_{k}$ is the dimensionality of the keys.
Each element in the output sequence of a Transformer layer is computed as a weighted sum of all values, with the weights assigned according to the dot products of the query with all keys. This allows each output element to dynamically attend to all input elements, facilitating better learning of contextual relationships in the data.
The Transformer uses multi-head attention to extend this mechanism:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O}$ (3)
$\text{where head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (4)
Here, $W^{O},W_{i}^{Q},W_{i}^{K},\text{and }W_{i}^{V}$ are parameter matrices, and $h$ is the number of attention heads. This design allows the model to jointly attend to information from different representation subspaces at different positions, enhancing the model’s ability to learn complex patterns.
Transformers are structured into an encoder-decoder architecture. The encoder maps an input sequence to a continuous representation that holds all the learned information of that sequence. The decoder then generates an output sequence from this representation. The entire process is facilitated by residual connections and layer normalization, enhancing training stability and model performance.
@@ III-C Black-Litterman Framework
The Black-Litterman (BL) model refines traditional portfolio optimization by blending market data with investor views. This integration improves investment strategies by aligning market conditions with personal risk preferences. Below are the key formulas that illustrate this method’s approach to asset allocation.
Formula (5) represents the calculation of the equilibrium return vector, $\Pi$, where $\delta$ denotes the risk aversion coefficient, $\Sigma$ is the covariance matrix of asset returns, and $w_{marker}$ are the market capital weights.
@@ Acknowledgment
This work is supported by the Guangdong Basic and Applied Basic Research Foundation (No. 2023B1515130002).